OpenStudy (anonymous):

Match the sequence (term) with the correct type of sequence (definition). Match Term Definition 128, 32, 8, 2, ... A) Arithmetic, common difference is 5 1, 3, 9, 27, ... B) Arithmetic, common difference is -3 5, 10, 15, 20, ... C) Geometric, common ratio is 0.25 20, 17, 14, 11, ... D) Geometric, common ratio is 3

OpenStudy (anonymous):

First one, you divide by 4 every time, so it would be C Second one, multiplied by 3, so it's D Third one, you add 5, so it's A Fourth one, you subtract 3, so it's B

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

can you help me with one more ill fan you

OpenStudy (anonymous):

Go ahead

OpenStudy (anonymous):

Jill has been watching the squirrels that live in her neighborhood. The number of squirrels changes each week. n f(n) 1 8 2 12 3 18 4 27 Which function best shows the relationship between n and f(n)? f(n) = 12(1.5)n f(n) = 12(1.5)n-1 f(n) = 8(1.5)n f(n) = 8(1.5)n-1

OpenStudy (anonymous):

i got D not sure if thats right

OpenStudy (anonymous):

The order of operations on it looks iffy. Are you sure the parenthesis are in the correct spots? With that being said I'd probably go with D.

OpenStudy (anonymous):

for the matching they said it was wrong

OpenStudy (anonymous):

THE MATCHING IS INCORRECT